OpenStudy (anonymous):

Solve for the variable: (b/2)+4=5

OpenStudy (judygreeneyes):

b + 8 = 10 b = 2

OpenStudy (anonymous):

Solve for b: (2b)/5+b/10-4 =6+3b-15